# Fast and Robust Visuomotor Riemannian Flow Matching Policy

### v1

@@ V-A3 Evaluation methodology
We evaluate RFMP, SRFMP, and DP using three key metrics: (1) The performance, computed as the average task-depending score across all trials, with $50$ trials for each simulated task, and $10$ trials for each real-world task; (2) The number of training epochs; and (3) The inference time. To provide a consistent measure of inference time across RFMP, SRFMP, and DP, we report it in terms of the number of function evaluations (NFE), which is proportional to the inference process time. Given that each function evaluation takes approximately the same time across all methods, inference time comparisons can be made directly based on NFE. For example, in real-world tasks, with an NFE of $2$, DP requires around $0.0075$\mathrm{s}$$, while RFMP and SRFMP take approximately $0.0108$\mathrm{s}$$ and $0.0113$\mathrm{s}$$, respectively. When NFE is increased to $5$, DP takes around $0.0175$\mathrm{s}$$, while RFMP and SRFMP require about $0.021$\mathrm{s}$$ and $0.0212$\mathrm{s}$$.
@@ V-B Push-T Tasks
Figure 4: Simulation benchmarks: Euclidean Push-T [2], Sphere Push-T, and four tasks from Robomimic [10]: Lift, Can, Square, Tool Hang.
@@ V-B1 Euclidean Push-T
First, we evaluate the performance of RFMP and SRFMP in the Euclidean case for different number of function evaluations in the testing phase. The models are trained with the default parameters described in Section V-A1. Table II shows the success rate of RFMP and SRFMP for different NFEs. We observe that both RFMP and SRFMP achieve similar success rates overall. While SRFMP demonstrates superior performance with a single NFE, RFMP achieves higher success rates with more NFEs. We hypothesize that this behavior arises from the fact that, due to the equality $\lambda_{\bm{x}}=\lambda_{\tau}$, the SRFMP conditional probability path resembles the optimal transport map between the prior and target distributions as in [15]. This, along with the stability framework of SRFMP, allows us to automatically choose the timestep during inference via (24), which enhances convergence in a single step. When comparing our approaches with DP, we observe that DP performs drastically worse than both RFMP and SRFMP for a single NFE, achieving a score of only $10.9\%$. Nevertheless, the performance of DP improves when increasing the NFE and matches that of our approaches for $10$ NFE.
TABLE II: Euclidean Push-T: Impact of NFE on policies
@@
Next, we ablate the action prediction horizon $T_{p}$, observation horizon $T_{o}$, learning rate $\eta$, and weight decay $w_{d}$ for RFMP and SRFMP. We consider $3$ different values for each, while setting the other hyperparameters to their default values, and test the resulting models with $5$ different NFE. For SRFMP, we additionally ablate the parameters $\lambda_{\bm{x}}$ and $\lambda_{\tau}$ for $3$ different ratios $\lambda_{\bm{x}}/\lambda_{\tau}$ and $4$ values for each ratio. Each setup is tested with $50$ seeds, resulting in a total of $2250$ and $4000$ experiments for RFMP and SRFMP. The results are reported in Tables III and IV, respectively. We observe that a short observation horizon $T_{o}=2$ leads to the best performance for both models. This is consistent with the task, as the current and previous images accurately provide the required information for the next pushing action, while the actions associated with past images rapidly become outdated. Moreover, we observe that an action prediction horizon $T_{p}=16$ leads to the highest score. We hypothesize that this horizon allows the model to maintain temporal consistency, while providing frequent enough updates of the actions according to the current observations. Concerning SRFMP, we find that $\lambda_{\bm{x}}=\lambda_{\tau}=2.5$ leads to the highest success rates. Interestingly, this choice leads to the ratio $\lambda_{\bm{x}}/\lambda_{\tau}=1$, in which case the flow of ${\bm{x}}$ follows the Gaussian CFM (6) of [15] with $\sigma\to 0$ for $\tau=[0,1]$, see [20, Cor 4.12]. In the next experiments, we use the default parameters resulting from our ablations, i.e., $T_{p}=16$, $T_{o}=2$, $\eta=1\times 10^{-4}$, $w_{d}=0.001$, and $\lambda_{\bm{x}}=\lambda_{\tau}=2.5$.
TABLE III: RFMP Hyperparameters Ablation on Euclidean Push-T
@@
TABLE IV: SRFMP Hyperparameters Ablation on Euclidean Push-T
@@ V-B2 Sphere Push-T
Next, we test the ability of RFMP and SRFMP to generate motions on non-Euclidean manifolds with the Sphere Push-T task. We evaluate two types of Riemannian prior distributions for RFMP and SRFMP, namely a spherical uniform distribution and wrapped Gaussian distribution. We additionally consider a Euclidean Gaussian distribution for DP. Notice that the actions generated by DP are normalized in a post-processing step to ensure that they belong to the sphere. The corresponding performance are reported in Table V. Our results indicate that the choice of prior distribution significantly impacts the performance of both RFMP and SRFMP. Specifically, we observe that RFMP and SRFMP with a uniform sphere distribution consistently outperform their counterparts with wrapped Gaussian distribution. We hypothesize that RFMP or SRFMP benefit from having samples that are close to the data support, which leads to simpler vector fields to learn. In other words, uniform distribution provides more samples around the data distribution, which potentially lead to simpler vector fields. DP exhibit poor performance with sphere-based prior distributions, suggesting its ineffectiveness in handling such priors. Instead, DP’s performance drastically improves when using a Euclidean Gaussian distribution and higher NFE. Note that this high performance does not scale to higher dimensional settings as already evident in the real-world experiments reported in Section V-D, where the effect of ignoring the geometry of the parameters exacerbates, which is a known issue when naively operating with Riemannian data [63]. Importantly, SRFMP is consistently more robust to NFE and achieves high performance with a single NFE, leading to shorter inference times for similar performance compared to RFMP and DP.
TABLE V: Impact of the prior distribution on Sphere Push-T Task
@@
TABLE VI: Influence of Integration Time on RFMP and SRFMP
@@ V-B3 Influence of Integration Time Boundary
We further assess the robustness of SRFMP to varying time boundaries on the Push-T tasks by increasing the time boundary during inference. The performance of both RFMP and SRFMP is summarized in Table VI with result presented for $\text{NFE}=3$ under the time boundaries $t=1$ and $t=1.2$, as well as for $\text{NFE}=4$ under the time boundaries $t=1.6$. Our results show that the performance of RFMP is highly sensitive to the time boundary, gradually declining as the boundary increases. In contrast, SRFMP demonstrates remarkable robustness, with minimal variation across different time boundaries. As illustrated in Figure 5, the quality of action series generated by RFMP noticeably deteriorates with increasing time boundaries, whereas SRFMP consistently delivers high-quality action series regardless of the time boundary.
V-C Simulated Robotic Experiments
Next, we evaluate RFMP and SRFMP on the well-known Robomimic robotic manipulation benchmark [10]. This benchmark consists of five tasks with varying difficulty levels. The benchmark provides two types of demonstrations, namely proficient human (PH) high-quality teleoperated demonstrations, and mixed human (MH) demonstrations. Each demonstration contains multi-modal observations, including state information, images, and depth data. We report results on four tasks (Lift, Can, Square, and Tool Hang) from the Robomimic dataset with $200$ PH demonstrations for training for both state- and vision-based observations. Note that the difficulty of the selected tasks becomes progressively more challenging. The score of each of the $50$ trials is determined by whether the task is completed successfully after a given number of steps ($300$ for Lift, $500$ for Can and Square, and $700$ for Tool Hang. The performance is then the percentage of successful trials.
Figure 6: Success rate on Robomimic tasks with state-based observations at different checkpoints. The models performance of Lift, Can, and Square tasks is checked every $10$ epochs throughout the $50$-epoch training process using $3$ NFE. For the Tool Hang task, the models are trained over $100$ epochs and checked every $20$ epochs using $10$ NFE.
@@ V-C1 State-based Observations
We first assess the training efficiency of RFMP and SRFMP and compare it against DP by analyzing their performance at different training stages. Figure 6 shows the success rate of the three policies as a function of the number of training epochs for each task. All policies are evaluated with $3$ NFE for Lift, Can, and Square, and with $10$ NFE for Tool Hang. We observe that RFMP and SRFMP consistently outperform DP across all tasks, requiring fewer training epochs to achieve comparable or superior performance. For the easier tasks (Lift and Can), both RFMP and SRFMP achieve high performance after just $20$ training epochs, while the success rate of DP remains low after $50$ epochs. This trend persists in the harder tasks (Square and Tool Hang), with RFMP and SRFMP reaching high success rates significantly faster than DP.
@@
TABLE VII: Success rate as a function of different NFE values on the state-based Robomimic tasks.
Task Lift Can Square Tool Hang

### v2

@@ V-A3 Evaluation methodology
We evaluate all policies using three key metrics: (1) The performance, computed as the average task-depending score across all trials, with $50$ trials for each simulated task, and $10$ trials for each real-world task; (2) The number of training epochs; and (3) The inference time. To provide a consistent measure of inference time across RFMP, SRFMP, and DP, we report it in terms of the number of function evaluations (NFE), which is proportional to the inference process time. Given that each function evaluation takes approximately the same time across all methods, inference time comparisons can be made directly based on NFE. For example, in real-world tasks, with an NFE of $2$, DP requires around $0.0075$\mathrm{s}$$, while RFMP and SRFMP take approximately $0.0108$\mathrm{s}$$ and $0.0113$\mathrm{s}$$, respectively. When NFE is increased to $5$, DP takes around $0.0175$\mathrm{s}$$, while RFMP and SRFMP require about $0.021$\mathrm{s}$$ and $0.0212$\mathrm{s}$$.
@@ V-B Push-T Tasks
Figure 6: Simulation benchmarks. Top row: Euclidean Push-T [2], Sphere Push-T, and five Robomimic tasks [10]: Lift, Can, Square, Tool Hang, Transport. Bottom row: Tasks of the Franka kitchen benchmarks, i.e., open microwave, put kettle on top burner, switch on the light, slide cabinet.
@@ V-B1 Euclidean Push-T
First, we evaluate the performance of RFMP and SRFMP in the Euclidean case for different number of function evaluations in the testing phase. The models are trained with the default parameters described in Section V-A1. Table III shows the success rate of RFMP and SRFMP for different NFEs. We observe that both RFMP and SRFMP achieve similar success rates overall. While SRFMP demonstrates superior performance with a single NFE, RFMP achieves higher success rates with more NFEs. We hypothesize that this behavior arises from the fact that, due to the equality $\lambda_{\bm{x}}=\lambda_{\tau}$, the SRFMP conditional probability path resembles the optimal transport map between the prior and target distributions as in [15]. This, along with the stability framework of SRFMP, allows us to automatically choose the timestep during inference via (24), which enhances convergence in a single step. When comparing our approaches with DP, we observe that DP performs drastically worse than both RFMP and SRFMP for a single NFE, achieving a score of only $10.9\%$. Nevertheless, the performance of DP improves when increasing the NFE and matches that of our approaches for $10$ NFE.
TABLE III: Euclidean Push-T: Impact of NFE on policies
@@
Next, we ablate the action prediction horizon $T_{p}$, observation horizon $T_{o}$, learning rate $\eta$, and weight decay $w_{d}$ for RFMP and SRFMP. We consider $3$ different values for each, while setting the other hyperparameters to their default values, and test the resulting models with $5$ different NFE. For SRFMP, we additionally ablate the parameters $\lambda_{\bm{x}}$ and $\lambda_{\tau}$ for $3$ different ratios $\lambda_{\bm{x}}/\lambda_{\tau}$ and $4$ values for each ratio. Each setup is tested with $50$ seeds, resulting in a total of $2250$ and $4000$ experiments for RFMP and SRFMP. The results are reported in Tables IV and V, respectively. We observe that a short observation horizon $T_{o}=2$ leads to the best performance for both models. This is consistent with the task, as the current and previous images accurately provide the required information for the next pushing action, while the actions associated with past images rapidly become outdated. Moreover, we observe that an action prediction horizon $T_{p}=16$ leads to the highest score. We hypothesize that this horizon allows the model to maintain temporal consistency, while providing frequent enough updates of the actions according to the current observations. Concerning SRFMP, we find that $\lambda_{\bm{x}}=\lambda_{\tau}=2.5$ leads to the highest success rates. Interestingly, this choice leads to the ratio $\lambda_{\bm{x}}/\lambda_{\tau}=1$, in which case the flow of ${\bm{x}}$ follows the Gaussian CFM (6) of [15] with $\sigma\to 0$ for $\tau=[0,1]$, see [20, Cor 4.12]. In the next experiments, we use the default parameters resulting from our ablations, i.e., $T_{p}=16$, $T_{o}=2$, $\eta=1\times 10^{-4}$, $w_{d}=0.001$, and $\lambda_{\bm{x}}=\lambda_{\tau}=2.5$.
TABLE IV: RFMP Hyperparameters Ablation on Euclidean Push-T
@@
TABLE V: SRFMP Hyperparameters Ablation on Euclidean Push-T
@@ V-B2 Sphere Push-T
Next, we test the ability of RFMP and SRFMP to generate motions on non-Euclidean manifolds with the Sphere Push-T task. We evaluate two types of Riemannian prior distributions for RFMP and SRFMP, namely a spherical uniform distribution and wrapped Gaussian distribution. We additionally consider a Euclidean Gaussian distribution for DP. Notice that the actions generated by DP are normalized in a post-processing step to ensure that they belong to the sphere. The corresponding performance are reported in Table VI. Our results indicate that the choice of prior distribution significantly impacts the performance of both RFMP and SRFMP. Specifically, we observe that RFMP and SRFMP with a uniform sphere distribution consistently outperform their counterparts with wrapped Gaussian distribution. We hypothesize that RFMP or SRFMP benefit from having samples that are close to the data support, which leads to simpler vector fields to learn. In other words, uniform distribution provides more samples around the data distribution, which potentially lead to simpler vector fields. DP exhibit poor performance with sphere-based prior distributions, suggesting its ineffectiveness in handling such priors. Instead, DP’s performance drastically improves when using a Euclidean Gaussian distribution and higher NFE. Note that this high performance does not scale to higher dimensional settings as already evident in the real-world experiments reported in Section V-E, where the effect of ignoring the geometry of the parameters exacerbates, which is a known issue when naively operating with Riemannian data [66]. Importantly, SRFMP is consistently more robust to NFE and achieves high performance with a single NFE, leading to shorter inference times for similar performance compared to RFMP and DP.
TABLE VI: Impact of the prior distribution on Sphere Push-T Task
@@
TABLE VII: Influence of Integration Time on RFMP and SRFMP
@@ V-B3 Influence of Integration Time Boundary
We further assess the robustness of SRFMP to varying time boundaries on the Push-T tasks by increasing the time boundary during inference. The performance of both RFMP and SRFMP is summarized in Table VII with result presented for $\text{NFE}=3$ under the time boundaries $t=1$ and $t=1.2$, as well as for $\text{NFE}=4$ under the time boundaries $t=1.6$. Our results show that the performance of RFMP is highly sensitive to the time boundary, gradually declining as the boundary increases. In contrast, SRFMP demonstrates remarkable robustness, with minimal variation across different time boundaries. As illustrated in Figure 7, the quality of action series generated by RFMP noticeably deteriorates with increasing time boundaries, whereas SRFMP consistently delivers high-quality action series regardless of the time boundary.
V-C Robomimic Benchmark
Next, we evaluate RFMP and SRFMP on the well-known Robomimic robotic manipulation benchmark [10]. This benchmark consists of five tasks with varying difficulty levels. The benchmark provides two types of demonstrations, namely proficient human (PH) high-quality teleoperated demonstrations, and mixed human (MH) demonstrations. Each demonstration contains multi-modal observations, including state information, images, and depth data. We report results on five tasks (Lift, Can, Square, Tool Hang, and Transport) from the Robomimic dataset with $200$ PH demonstrations for training for both state- and vision-based observations. Note that the difficulty of the selected tasks becomes progressively more challenging. The score of each of the $50$ trials is determined by whether the task is completed successfully after a given number of steps ($300$ for Lift, $500$ for Can and Square, $700$ for Tool Hang, and $500$ for Transport. The performance is then the percentage of successful trials.
Figure 8: Success rate (mean and standard deviation) on Robomimic tasks with state-based observations at different checkpoints. The models performance of Lift, Can, and Square tasks is checked every $10$ epochs throughout the $50$-epoch training process using $3$ NFE. For the Tool Hang task, the models are trained over $100$ epochs and checked every $20$ epochs using $10$ NFE.
@@ V-C1 State-based Observations
We first assess the training efficiency of RFMP and SRFMP and compare it against DP by analyzing their performance at different training stages. Figure 8 shows the success rate of the three policies as a function of the number of training epochs for the tasks Lift, Can, Square, and Tool Hang. All policies are evaluated with $3$ NFE for Lift, Can, and Square, and with $10$ NFE for Tool Hang. We observe that RFMP and SRFMP consistently outperform DP across all tasks, requiring fewer training epochs to achieve comparable or superior performance. For the easier tasks (Lift and Can), both RFMP and SRFMP achieve high performance after just $20$ training epochs, while the success rate of DP remains low after $50$ epochs. This trend persists in the harder tasks (Square and Tool Hang), with RFMP and SRFMP reaching high success rates significantly faster than DP.
@@
TABLE VIII: Success rate as a function of different NFE values on the state-based Robomimic tasks.
Lift
NFE $1$ $2$ $3$ $5$ $10$
RFMP ${0.992\pm 0.010}$ ${\bm{0.992\pm 0.010}}$ ${\bm{0.992\pm 0.010}}$ ${0.992\pm 0.010}$ ${0.992\pm 0.010}$
SRFMP ${\bm{1.000\pm 0.000}}$ ${\bm{0.992\pm 0.010}}$ ${\bm{0.992\pm 0.010}}$ ${\bm{0.996\pm 0.008}}$ ${\bm{1.000\pm 0.000}}$
DP ${0.008\pm 0.010}$ ${0.756\pm 0.034}$ ${0.948\pm 0.016}$ ${0.956\pm 0.015}$ ${0.976\pm 0.015}$
Can
NFE $1$ $2$ $3$ $5$ $10$
RFMP ${0.976\pm 0.023}$ ${\bm{0.996\pm 0.008}}$ ${\bm{0.996\pm 0.008}}$ ${\bm{0.996\pm 0.008}}$ ${\bm{0.996\pm 0.008}}$
SRFMP ${\bm{0.980\pm 0.022}}$ ${0.980\pm 0.013}$ ${\bm{0.996\pm 0.008}}$ ${0.992\pm 0.010}$ ${0.968\pm 0.020}$
DP ${0.004\pm 0.008}$ ${0.340\pm 0.013}$ ${0.836\pm 0.020}$ ${0.924\pm 0.015}$ ${0.908\pm 0.010}$
Square
NFE $1$ $2$ $3$ $5$ $10$
RFMP ${\bm{0.792\pm 0.027}}$ ${\bm{0.848\pm 0.020}}$ ${\bm{0.920\pm 0.018}}$ ${\bm{0.896\pm 0.041}}$ ${\bm{0.912\pm 0.016}}$
SRFMP ${0.776\pm 0.029}$ ${0.800\pm 0.052}$ ${0.828\pm 0.016}$ ${0.824\pm 0.008}$ ${0.848\pm 0.030}$
DP ${0.012\pm 0.010}$ ${0.384\pm 0.023}$ ${0.628\pm 0.016}$ ${0.672\pm 0.020}$ ${0.684\pm 0.015}$
Tool Hang
NFE $1$ $2$ $3$ $5$ $10$
RFMP ${0.152\pm 0.020}$ ${\bm{0.316\pm 0.023}}$ ${\bm{0.368\pm 0.032}}$ ${\bm{0.572\pm 0.027}}$ ${\bm{0.716\pm 0.023}}$
SRFMP ${\bm{0.240\pm 0.028}}$ ${0.224\pm 0.020}$ ${0.308\pm 0.020}$ ${0.516\pm 0.023}$ ${0.568\pm 0.035}$
DP ${0.000\pm 0.000}$ ${0.008\pm 0.010}$ ${0.008\pm 0.010}$ ${0.092\pm 0.024}$ ${0.092\pm 0.016}$
TABLE IX: Jerkiness of predicted robot trajectories for different NFE on the robomimic tasks with state-based observations. All values are expressed in thousands, where the lower the smoother the prediction.
Lift
NFE $1$ $2$ $3$ $5$ $10$
RFMP ${\bm{9.97\pm 1.23}}$ ${\bm{8.80\pm 0.19}}$ ${9.06\pm 0.08}$ ${8.96\pm 0.08}$ ${8.49\pm 0.21}$
SRFMP ${10.32\pm 0.41}$ ${9.90\pm 031}$ ${\bm{8.00\pm 0.14}}$ ${9.39\pm 0.23}$ ${9.03\pm 0.19}$
DP ${329.20\pm 47.44}$ ${13.96\pm 0.89}$ ${8.06\pm 0.42}$ ${\bm{7.48\pm 0.13}}$ ${\bm{5.9\pm 0.23}}$
Can
NFE $1$ $2$ $3$ $5$ $10$
RFMP ${7.88\pm 0.14}$ ${\bm{6.38\pm 0.40}}$ ${\bm{6.37\pm 0.17}}$ ${6.40\pm 0.07}$ ${7.10\pm 1.30}$
SRFMP ${\bm{7.64\pm 0.23}}$ ${7.63\pm 0.18}$ ${6.69\pm 0.27}$ ${6.98\pm 0.13}$ ${6.94\pm 0.08}$
DP ${526.00\pm 65.50}$ ${18.98\pm 2.82}$ ${6.68\pm 0.25}$ ${\bm{6.28\pm 0.35}}$ ${\bm{6.22\pm 0.16}}$
Square
NFE $1$ $2$ $3$ $5$ $10$
RFMP ${\bm{9.14\pm 0.19}}$ ${\bm{6.22\pm 0.29}}$ ${\bm{5.54\pm 0.24}}$ ${\bm{5.17\pm 0.16}}$ ${\bm{4.36\pm 0.30}}$
SRFMP ${9.16\pm 0.31}$ ${9.43\pm 0.16}$ ${7.28\pm 0.97}$ ${7.70\pm 0.50}$ ${7.74\pm 0.26}$
DP ${548.20\pm 54.77}$ ${20.26\pm 2.47}$ ${7.24\pm 0.54}$ ${10.02\pm 2.64}$ ${7.62\pm 1.44}$
Tool Hang
NFE $1$ $2$ $3$ $5$ $10$
RFMP ${\bm{4.49\pm 0.35}}$ ${\bm{4.26\pm 0.19}}$ ${\bm{4.65\pm 0.32}}$ ${\bm{5.07\pm 0.19}}$ ${\bm{4.66\pm 0.31}}$
SRFMP ${5.24\pm 0.18}$ ${5.10\pm 0.20}$ ${5.23\pm 0.38}$ ${5.62\pm 0.42}$ ${5.62\pm 0.40}$
DP ${699.40\pm 90.91}$ ${9.52\pm 0.30}$ ${8.79\pm 1.51}$ ${7.12\pm 1.94}$ ${6.96\pm 1.35}$
Next, we evaluate the performance of the policies for different NFE in the testing phase. For RFMP and SRFMP, we use the $50$-epoch models for Lift, Can, and Square, and the $100$-epoch model for Tool Hang. DP is further trained for a total of $300$ epochs and we select the model at the best validation epoch. The results are reported in Table VIII. RFMP and SRFMP outperform DP for all task and all NFE, even though DP was trained for more epochs. Moreover, we observe that RFMP and SRFMP are generally more robust to low NFE than DP. They achieve $100$% success rate at almost all NFE values for the easier Lift and Can tasks, while DP’s performance drastically drops for $1$ and $2$ NFE. We observe a similar trend for the Square task, where the performance of RFMP and SRFMP slightly improves when increasing the NFE. The performance of all models drops for Tool Hang, which is the most complex of the four considered tasks. In this case, the performance of RFMP and SRFMP is limited for low NFE values and improves for higher NFE. DP performs poorly for all considered NFE values. Table IX reports the jerkiness as a measure of the smoothness of the trajectories generated by the different policies. We observe that RFMP and SRFMP produce arguably smoother trajectories than DP for low NFE, as indicated by the lower jerkiness values. The smoothness of the trajectories becomes comparable for higher NFE. In summary, both RFMP and SRFMP achieve high success rates and smooth action predictions with low NFE, enabling faster inference without compromising task completion.
TABLE X: Training time (in seconds) per epoch on Robomimic vision-basedLift and Square.
Task RFMP SRFMP DP CP
Lift ${17.49\pm 0.03}$ ${17.63\pm 0.15}$ ${16.07\pm 0.20}$ ${31.44\pm 0.27}$
Square ${57.62\pm 0.30}$ ${56.67\pm 0.18}$ ${53.16\pm 0.16}$ ${102.86\pm 0.61}$
V-C2 Vision-based Observations
Next, we assess our models performance when the vector field is conditioned on visual observations. We consider the tasks Lift, Can, Square, and Transport with the same policy settings and networks (see Table I) [^1] Note that we omit Tool Hang due to its significant computational cost.. Each observation ${\bm{o}}^{s}$ corresponds to a visual embedding derived from camera images. For Lift, Can, and Square, we use one over-the-shoulder and one in-hand camera. For Transport, which involves bimanual manipulation, we use two in-hand cameras and two over-the-shoulder cameras. We train the models for a total $100$ epochs for Lift, Can, and Square and for $200$ epochs for Transport and use the best-performing checkpoint for evaluation.
The performance of different policies is reported in Table XI. RFMP and SRFMP consistently outperform DP and CP on all tasks, regardless of the NFE. As for the previous experiments, our models are remarkably robust to changes in NFE compared to DP. Importantly, SRFMP consistently outperforms RFMP for $1$ and $2$ NFE. Regarding Can and Square tasks, SRFMP with $1$ NFE achieved performance on par with RFMP using $3$ NFE. This efficiency gain showcases the benefits of enhancing the policies with stability to the target distribution for reducing their inference time.
In contrast to CP, RFMP and SRFMP rely on a simple, single-stage training pipeline, thus featuring easier and faster training in addition to fast inference. The training time per epoch of each policy on task Lift and Square tasks is summarized in Table X. All experiments were conducted on an NVIDIA RTX 4060Ti GPU using the same batch size across policies. The result indicate that RFMP, SRFMP, and DP have comparable per-epoch training times. In contrast, CP requires approximately twice as much time per epoch, primarily due to its training procedure involving distillation from a teacher policy. When accounting for the additional time required to pretrain DP the total training time per epoch for CP becomes roughly three times that of the other methods.
TABLE XI: Success rate as a function of NFE on vision-based robomimic tasks.
Task Lift Can Square Transport